# A Survey on Heterogeneous Computing Using SmartNICs and Emerging Data Processing Units (Expanded Preprint)

### v1

@@ 6.1 Broad Studies and Benchmarks
Benchmarking these devices is essential for assessing their performance and efficiency in different applications. Such benchmarks may measure factors such as CPU or network throughput, latency, and energy efficiency, providing insights into the DPU’s capabilities under various workloads. These benchmarks help in optimizing hardware and software configurations for specific use cases, ensuring they deliver optimal performance in their target applications. While a few of the following works are broad studies, and others are in-depth general device benchmarks, other works only measure their specific contribution’s effect; since most works do this to some extent, we will identify only a few contribution-specific works of note here. Table 4 includes a mixture of selections of both.
Note that, due to the relatively small number of works performing thorough benchmarks, we have not specified which SmartNIC type each work is using, but instead have categorized by benchmark topic.
@@
Topic or Benchmark Category General Contribution-Specific
Works Creating Benchmark Suites [63, 91]
Non-Benchmark Studies [49, 65, 19, 12, 29, 5, 24, 62, 75, 96, 34]
Whitepapers & Specs [21, 11, 67, 36, 61, 96, 14, 1]
Compute Power [81, 59, 60] [40]
Communication (Throughput, Latency) [49, 81, 92, 59, 60, 43, 24, 11] [73, 43, 98, 51, 25, 72, 40, 30, 74, 53, 26]
Memory Access (Local, Distributed) [81, 59, 60] [102]
Storage (Local, Distributed) [59, 11] [92, 104, 77, 45, 17, 18, 50, 73, 33, 26]
Compression [49, 58, 55, 56, 59] [54, 70, 17, 18]
Other ASIC Accelerators [81, 59]
Applications, System Services [11] [32, 43, 8, 60, 85, 45, 57, 83, 46, 13]
Engergy Savings [21] [69, 105, 100, 84]
Scalability [43, 23, 32, 31, 27, 37, 38]
Broad Studies and Advice: In [81], Sun et al. use deeper evaluations of compute power and communication overhead, performed on a BlueField, to recommend four experiment-supported pieces of advice for using off-path SmartNICs to improve performance:
@@
Finally, they discuss possible refutations of performance gains shown by a few other papers, demonstrating the difficulty of creating any effective offloads for off-path SmartNICs. Xingda Wei et al. [92] also perform a comprehensive study of the Bluefield-2, focusing on file system and key-value store (KV-store) communication paths, and using multiple paths concurrently, to discover substantial performance improvements. They then give advice on finding such potential optimizations, which they, too, have validated through their studies. Jianshen Liu et al. [59] also evaluate the capabilities of the network and computing aspects of the NVIDIA BlueField-2, and showed that while the device is flexible, it is easy to bottleneck and overwhelm it. They recommend a selection of operations for offloading, which this DPU performs particularly well at.
A Big-Picture View on Cloud Architecture: Some years back, Caulfield et al. [12] published an invited work which is not so much a broad study or benchmark, as it is an analysis of the current state of the research regarding fully programmable networks. They provide an in-depth discussion towards applying FPGA SmartNICs to create a more flexible cloud infrastructure, going beyond the use of SmartNICs for offloading specific tasks, and discussing the architecting of highly dynamic data center systems to meet the growing needs for abstractions to power infrastructure, applications, and developers alike. We find this discussion interesting, as it appears to reflect how network hardware and software have developed since its publication; for example, lacking reference to proper DPUs, which now possess some of the abstractions they call for. The reader may find its big-picture view enlightening.
In a more recent review, Ryan E. Grant et al. [29] discuss the design of SmartNICs and how they can be used to offload current runtime software, as well as enable future in-network runtime systems. Döring et al. [19] discuss the demand for these devices, as well as their creation and use in attempting to meet the growing needs surrounding Software-Defined Networking (SDN). They highlight how SmartNICs can be used to tackle the lack of network flexibility, making virtualized networks more feasible. The paper also aims to provide a universal definition for the term “SmartNIC”. Luca Barsellotti et al. [5] go on to discuss three different use cases for DPUs at the edge: network monitoring for 5G networks, power efficiency in edge-to-cloud, and added network security. Horany et al. [34] describe a recommended model for offloading and allocating SmartNIC resources across a network, using gRPC, among other tools.
On a related note, Justine Sherry [75] makes an argument that instead of considering DPUs only for compute acceleration and offload, these devices can develop to become the overall data movement controller (DMC) for a compute node with a multi-core CPU and other accelerators like GPUs. This is a new way to see the role of emerging DPUs, when compared to traditionally making the host responsible for both computations and orchestrating data movement. The paper presents challenges and open questions in realizing this vision in the data center environment.
@@
Finally, recent detailed surveys can be very instructive, such as those presented by Kfoury et al. [49] for general information about SmartNICs, and Nickel et al. [65] for information on in-network computing with SmartNICs.
@@
Compression Studies: Recently, there have been a number of works studying, more deeply, the benefits of using DPUs in order to accelerate data compression tasks, utilizing the compression related hardware building blocks in NVidia’s Bluefield SmartNICs. One particularly relevant work is that of Yuke Li et al. [55, 56], who first examine the performance of lossy and lossless compression on NVIDIA BlueField DPUs (particularly SZ3, DEFLATE, and zlib) using seven real-world datasets, underlining the potential of offloading compression tasks from host CPUs to DPUs to enhance the performance in data-intensive applications. Later, they move to accelerate such compression tasks, which we will describe in Section 6.2, along with a number of other works on compression.
@@
Creating New Benchmark Suites: And finally, where there are those using benchmarks, there must also be those creating useful new benchmarks. Wang et al. [91] proposed ‘DPUBench’ an application-driven benchmark suite, covering three common categories of DPU applications: network, storage, and security. The suite includes a scalable framework for benchmarking how well DPUs handle tasks relating to those categories. Similarly, Michalowicz et al. [63] created a micro-benchmark suite, also called ‘DPU-Bench’, for helping HPC researchers to determine how many processes can be offloaded to a DPU without degrading performance, given various factors.
6.2 DPUs in Commercial and Infrastructure Applications
The commercial application of DPUs has become increasingly prevalent in modern computing environments. Thus, much of the current research around DPUs is in using them to improve the infrastructure itself of the data centers they belong to, since their ability to accelerate network communication and offload tasks relating to packet processing, security, infrastructure processing, storage, virtualization, cloud computing etc. can provide an important advantage to the companies that use them. As these are the primary familiar and intended applications of SmartNICs, these areas of research tend to show many positive results, as described in the following papers.
@@ 6.2 DPUs in Commercial and Infrastructure Applications
However, as the line is often very blurred between these and more unconventional applications of SmartNICs, in this section we will also present those works which experiment with using SmartNICs for various other kinds of offloads and co-processing tasks, reaching far into the realm of applications as well as infrastructure. The flexibility of DPUs has led to their adoption in a surprising range of domains, beyond simply accelerating data center infrastructure and security; while they are quickly becoming crucial in that realm, their surprising application to unexpected problem domains has also skyrocketed. From blockchain technologies to molecular dynamics simulations [43, 86], they are being treated as co-processors, data management and analytics devices, machine learning accelerators, and more. Table 5 gives an attempted breakdown of the many overlapping topics within this category.
@@
Hardware Implementation, FPGAs [102, 53, 24, 10, 74] (See also Section 6.3)
Distributed Storage [102, 63, 28]
@@
Network Offloading and Load-Balancing Frameworks: One of the earlier pioneers in the area of SmartNIC packet-processing offloading is the work of Antoine Kaufmann et al. [45], a direct memory access (DMA) interface called ‘FlexNIC’. Since then, other works have continued to forge ahead. Ming Liu et al. [60] presented ‘iPipe’, an “actor-based” framework to aid in offloading varied, distributed applications to SmartNICs in a way that maximally uses the SmartNICs’ cores, despite differences in task execution costs. Evaluations of iPipe show significant core savings (up to 3.1 decent Intel cores in a 10Gbps scenario) and latency reductions by 23.0 $\mu$s when offloading to SmartNICs for data analytics, transaction systems, and KV-stores. However, [81] questions the efficacy of iPipe’s design for off-path SmartNICs. J. Ravi et al. [70] introduced the ‘Runway’ framework, which uses an object storage abstraction to allow for computation on data while it’s in-transit, and offloading compression to a heterogeneous device. This framework is specifically designed to be adaptable in order to execute user-defined functions during runtime. They perform tests on both CPUs, DPUs, and GPUs. Finally, Fuhrer et al. [25] use on-NIC reinforcement learning for network congestion control, but we will discuss this work in Section 6.3.
A potential way to use SmartNICs to improve disk and network I/O operations is by rearchitecting the TCP stack itself, as done in [51] by Taehyun Kim et al. They present ‘IO-TCP’, which improves content delivery by cleverly splitting its stack of operations between the host CPU and the SmartNIC. Since offloading the whole stack to the SmartNIC would leave data inaccessible to the CPU, and the SmartNIC isn’t fast enough to compete for some operations, they differentiate between “control plane” operations to be handled by the CPU, and “data plane” operations to be managed by the SmartNIC. This reduces cache pollution between operations with large amounts of disk I/O. Their work is tested on the Bluefield-2 DPU, and demonstrates approximately double the throughput in cases with high CPU congestion, even compared to systems using far more CPU cores.
@@
Hardware and FPGA-Based; Network, Packets, & Transactions: Some works, such as that of Jie Zhang et al. in [102], aim to improve the design of SmartNICs themselves. Their efforts demonstrate ‘SmartDS’, an FPGA prototype of a SmartNIC with the ability to split message headers and bodies between the host and SmartDS, improving throughput for distributed storage systems, with potential for accelerating and reducing cloud storage infrastructure by offloading compression, reducing middle-tier servers required for disagregated block storage, etc. They test their prototype in the middle-tier of a distributed system, and show up to 4.3x throughput compared to CPU-based middle-tier management, and also reducing the average latency by 2.6x. Similarly, Li et al. develop ‘AlNiCo’ [53], an on-NIC network transaction scheduling system which aims to offload the scheduling of transaction requests and responses to specific host CPUs, in a manner which reduces both overhead and resource contention, and which adapts to to feedback during dynamic workloads. Their implementation involves both hardware and software solutions, using an FPGA-based SmartNIC to program their own hardware acceleration which they feed feature vectors comprised of request information. Their solution also involves RDMA (Remote Direct Memory Access), and the development of their own “scheduling-enabled” RPC (Remote Procedure Call). They show improvements to both throughput and latency. Also, Shan Yizhou et al. [74] fill a niche by presenting the first SmartNIC that is both multi-tenant for virtualization needs, and programmable but hardware-based, called ‘SuperNIC’. They focused on optimizing execution of entire connected groups (flows or DAGs-Directed acyclic Graphs) of network tasks, and included various optimization techniques resulting in low overhead and high throughput compared to the baseline. Furthermore, the researchers proposed a fair sharing mechanism for hardware resources in their SuperNIC. As for traffic analysis, Siracusano et al. [27] use a hardware-implemented binary neural network, described in section 6.3), to speed things up by two orders of magnitude. Finally, Brunella et al.’s ‘Hyperion’ [10], described below, is a more direct example of modifying DPU hardware, in pursuit of better distributed storage systems.
[…]
@@
Distributed File Systems: Others have also been able to achieve substantial speedups by focusing on distributed file systems (DFS). LineFS, a persistent memory (PM) distributed file system (DFS), was first presented by Jongyul Kim et al. in [50], a paper describing its design, implementation, and evaluation. LineFS resides within the SmartNIC itself to offload DFS operations and reduce CPU contention via in-network processing, while still taking advantage of new client-local persistent memory (PM) to create speedups by preferring client-local behavior. They test their work on a Mellanox Bluefield and compare it to Assise (a current PM DFS), showing improvement in DFS availability, latency of LevelDB operations by up to 80%, and throughput in Filebench by as much as 79%. However, [81] notes a limited range of situations in which LineFS improves performance. Xingda Wei et al. [92] perform case studies using both LineFS (an on-SmartNIC DFS) and DrTM-KV (an RDMA-based disagregated KV-store), focusing on the management of such systems, and the idea of using multiple communication paths concurrently, achieving up to 30% and 25% improvements, respectively, using a Bluefield-2 DPU.
[…]
@@
The concept of a DPU has also grown in an interesting direction with Brunella et al.’s ‘Hyperion’, which replaces host CPU-based devices as the primary manager of distributed NVMe storage  [10], thus reimagining typical architectures to eliminate the need for many CPU server devices.
@@
Compression: Yuke Li et al. [54]. They introduced a library ‘PEDAL’ that optimizes data compression designs utilizing the DPU’s hardware capabilities. They also highlight the challenges of existing compression methods due to high computational demands, and present a co-design with the MPICH MPI library which achieves a speedup up to 101x in compression time, and reducing the communication latency by up to 88x, compared to not using ASICs. Lastly, they focused on enhancing the data compression in communication-oriented HPC scenarios. Chen Ding et al. [17, 18] also choose to offload compression to the DPU with D2Comp, in order to accelerate LSM-tree compaction; Ravi et al. [70] do so with Runway for in-transit computation; and Zhang et al. [102] with SmartDS (all detailed in other parts of this survey). Others, such as Jianshen [58], of course make use of the compression ASICs as well.
@@
KV-Stores, Databases, & Big Data Systems: In [57], Jiaxin Lin et al. focus their efforts on the shuffle processes (map-reduce, etc.) of data-intensive applications, and more generally on dynamically partially offloading operations requiring intermediate data exchange between compute nodes. They present ‘SmartShuffle’ for coordinating offload to maximize usage of both the CPU and SmartNIC simultaneously. They attach it to Spark and test it on a Broadcom Stingray, achieving up to 40% faster performance on TPC-H benchmarks than both Spark alone and Spark RDMA. Somewhat related, Henry N. Schuh et al. [73] introduce ‘Xenic’, a system that employs an asynchronous, aggregated execution paradigm with adaptable point-to-point communication between SmartNICs, improving both both network and core efficiency in performing sharded data store transactions. Their experimental findings demonstrate that Xenic achieves notable improvements on three different benchmarks, at least doubling throughput and significantly reducing latency. However, [81] questions Xenic’s use in off-path Smart-NICs.
@@
As for more specific aspects of such systems, Chen Ding et al. [17, 18] proposed another improvement, ‘DComp’ and then ‘D²Comp’, wherein they integrate DPU-offloaded LSM-tree compaction with RocksDB. This approach was evaluated using NVIDIA’s Bluefield-2 DPU, the results of which demonstrated the effectiveness of the DPU solution, accelerating compaction performance by up to 4x, write and read throughput by up to 3.2x and 1.4x respectively, and boasting a reduction in host CPU contention. Also, Haiyang Shi et al. [77] proposed a set of coherent in-network erasure coding EC primitives, named ‘INEC’ (building on their previous work named ‘TriEC’ [76]), to more easily allow offloading of EC to SmartNICs despite the variety of schemes EC follows. The researchers conducted experiments and demonstrated that NIC implementations of INEC primitives, for various EC schemes, with a KV store co-designed with INEC, results in a substantial improvement at the 50th, 95th, and 99th percentiles in end-to-end throughput, write performance, and degraded read performance, achieving enhancements of up to 99.57%, 47.30%, and 49.55% respectively Works belonging to other sections, such as [92], may also deal with KV-stores.
@@
Lastly, when dealing with network data flows through SmartNICs, Jianshen Liu et al. [58] recommend the use of Apache Arrow at the foundation, because its data format can save data transformation time. The authors share their experience in adjusting a partitioning algorithm for particle data to work with Apache Arrow.
@@
5G Network Infrastructure: Yan yan et al. [99] introduced a SmartNIC solution that is P4-enabled and FPGA-based. This solution was specifically developed to cater to the networking requirements of web-scale cloud and the 5G / beyond-5G era. The authors demonstrated the application of this solution in a 5G environment, with a particular focus on network slicing from edge data center to core data center. The experimental outcomes showed that the system can achieve a throughput of up to 84.8 Gbps, even using only a single CPU core. Additionally, it was observed that the incorporation of hardware-offloaded segment routing with P4 can achieve up to 30% higher bandwidth. Later, Justine Cris Borromeo et al. [8] add a 5G DU Low-PHY layer to an FPGA SmartNIC using the OpenCL framework. They demonstrate that incorporating these 5G functionalities into a SmartNIC as an offload, as opposed to other current solutions for a 5G network, can result in notable reductions in both processing time and power consumption.

### v2

@@ 5 Broad Studies and Benchmarking of DPUs
In this section, we identify those works which are broad studies, or which perform deeper benchmarks of DPUs. Benchmarking these devices is essential for assessing their performance and efficiency in different applications. Such benchmarks may measure factors such as CPU or network throughput, latency, and energy efficiency, providing insights into the DPU’s capabilities under various workloads. These benchmarks help in optimizing hardware and software configurations for specific use cases, ensuring that they deliver optimal performance in their target applications. While a few of the following works are broad studies, and others are in-depth general device benchmarks, other works only measure their specific contribution’s effect; since most works do this to some extent, we will identify only a few contribution-specific works of note here. Table 5 includes a mixture of selections of both.
Note that, due to the relatively small number of works performing thorough benchmarks, we have not specified in Table 5 which SmartNIC type each work is using, but instead have categorized them by benchmark topic. In this work we do not compile a comparison of benchmarks between available DPU devices, in part because the existing works tend to primarily provide information on only a few more popular models, such as NVIDIA BlueField DPUs; nor do we directly compare those works’ results, seeing that their methodology, specific tests or datasets, device or code configurations, and even the ASICS and other aspects of the devices being benchmarked differ so much as to make them incomparable in many cases. Instead, we refer the reader to the well-performed benchmarks described by the papers directly following.
Topic or Benchmark Category General Contribution-Specific
Works Creating Benchmark Suites [33, 25, 20, 51]
Non-Benchmark Studies [1, 78, 48, 9, 54, 19, 49, 79, 44, 59]
Whitepapers & Specs [2, 43, 28, 41, 36, 44, 38, 64, 35, 3, 16]
Compute Power [23, 46, 33, 24, 80] [45, 56, 74]
Communication (Throughput, Latency) [1, 23, 81, 46, 25, 33, 24, 30, 43] [37, 30, 61, 71, 82, 52, 45, 26, 83, 70, 69]
Memory Access (Local, Distributed) [23, 46, 25, 33, 24] [84]
Storage (Local, Distributed) [46, 25, 33, 43] [81, 85, 73, 86, 87, 88, 37, 68, 69]
Compression [1, 18, 89, 46, 33] [72, 47, 86, 87]
Other ASIC Accelerators [1, 23, 46, 33, 80]
Applications, System Services [43] [76, 30, 90, 24, 75, 91, 92, 58, 93, 56, 74]
Engergy Savings [2] [94, 77, 32, 95]
Scalability [30, 96, 76, 97, 98, 55]
@@
Broad Studies and Advice: Sun et al. [23] use deeper evaluations of compute power and communication overhead, performed on a BlueField, to recommend four experiment-supported pieces of advice for using off-path SmartNICs to improve performance:
@@
Finally, they discuss possible refutations of performance gains shown by a few other papers, demonstrating the difficulty of creating truly effective offloads for off-path SmartNICs. Wei et al. [81] also perform a comprehensive study of the BlueField-2, focusing on file system and key-value store (KV-store) communication paths, and using multiple paths concurrently, to discover notable performance improvements. They then give advice on finding such potential optimizations, which they have validated through their studies. J. Liu et al. [46] also evaluate the capabilities of the network and computing aspects of the NVIDIA BlueField-2, and showed that while the device is flexible, it is easy to bottleneck and overwhelm it. They recommend a selection of operations for offloading, which this DPU performs particularly well at. Kashyap et al. [25] focus on evaluating the DMA and RDMA paths of multi-generational DPUs, and usefully identify a number of device idiosyncrasies.
Chen et al. [80], instead of measuring the capabilities of the CPU, delve into the datapath accelerator (DPA) of the BlueField-3, which is a small real-time unit with low resources but a high thread count, primarily intended to accelerate the DPU’s NIC system. They find it can be efficiently applied to certain types of offloads, and give advice thereon.
@@
Finally, recent detailed surveys can be very instructive, such as those presented by Kfoury et al. [1] for general information about SmartNICs, and Nickel et al. [78] for information on in-network computing with SmartNICs. Works discussing benchmark suites may also give broad useful information, such as that given by ‘dpBento’ [33] regarding the types of tasks DPUs perform most effectively.
A Big-Picture View on Cloud Architecture: Some years back, Caulfield et al. [9] published an invited work which is not so much a broad study or benchmark, as it is an analysis of the current state of the research regarding fully programmable networks. They provide an in-depth discussion towards applying FPGA SmartNICs to create a more flexible cloud infrastructure, going beyond the use of SmartNICs for offloading specific tasks, and discussing the architecting of highly dynamic data center systems to meet the growing needs for abstractions able to power infrastructure, applications, and developers alike. We find this discussion interesting, as it appears to reflect how network hardware and software have developed since its publication; for example, they lack reference to proper DPUs, which now possess some of the abstractions the paper calls for. They provide an informative big-picture view.
In a more recent review, Grant et al. [54] discuss the design of SmartNICs and how they can be used to offload current runtime software, as well as enable future in-network runtime systems. Döring et al. [48] discuss the demand for these devices, as well as their creation and use in attempting to meet the growing needs surrounding Software-Defined Networking (SDN). They highlight how SmartNICs can be used to tackle the lack of network flexibility, making virtualized networks more feasible. The paper also aims to provide a universal definition for the term “SmartNIC”. Barsellotti et al. [19] go on to discuss three different use cases for DPUs at the edge: network monitoring for 5G networks, power efficiency in edge-to-cloud, and added network security. Horany et al. [59] recommended and describe a model for offloading and allocating SmartNIC resources across a network, using gRPC, among other tools.
On a related note, Justine Sherry [79] makes an argument that instead of considering DPUs only for compute acceleration and offload, these devices can develop to become the overall data movement controller (DMC) for a compute node with a multi-core CPU and other accelerators like GPUs. This is a new paradigm for the role of emerging DPUs, compared to traditionally making the host responsible for both computations and orchestrating data movement. The paper presents challenges and open questions in realizing this vision in the data center environment, including a list of unresolved queries regarding how the hardware and software systems of the future will adapt to accommodate such a dedicated ‘NIC-DMC’ operating independently from the CPU complex.
@@
Compression Studies: Recently, there have been a number of works studying, more deeply, the benefits of using DPUs in order to accelerate data compression tasks, utilizing the compression related hardware building blocks in NVIDIA’s BlueField SmartNICs. One particularly relevant work is that of Li et al. [89], who first examine the performance of lossy and lossless compression on NVIDIA BlueField DPUs (particularly SZ3, DEFLATE, and zlib) using seven real-world datasets, underlining the potential of offloading compression tasks from host CPUs to DPUs to enhance the performance in data-intensive applications. Later, they move to further accelerate such compression tasks, which we will describe in Section 6.1, along with a number of other works on compression.
@@
Creating New Benchmark Suites: And finally, where there are works using benchmarks, there are also works creating new benchmarks. ‘dpBento’ [33] is a benchmark suite specializing in data processing; its creators share helpful insights found by their own use of it. We find this paper to be of particular interest because of these insights; for example, they observe that some DPUs may outperform their hosts at processing smaller operands and floating-point operations, which may be of interest to those offloading some types of data-oriented tasks, such as the geospatial queries of Kaymak et al. [58], discussed later. Wang et al. [51] proposed ‘DPUBench’ an application-driven benchmark suite, covering three common categories of DPU applications: network, storage, and security. The suite includes a scalable framework for benchmarking how well DPUs handle tasks relating to those categories. Nearly of the same name, Michalowicz et al. [20] created a micro-benchmark suite, also called ‘DPU-Bench’, for helping HPC researchers to determine how many processes can be offloaded to a DPU without degrading performance, given various factors. Kashyap et al. [25] create a microbenchmark, DPUDMABench, for benchmarking DMA capabilities.
6 Application Areas of DPUs
The remainder of this survey might be considered a form of topical index, intending to point researchers to more complete treatments of each area than we can give here, categorizing and grouping the main body of papers we have reviewed by the various applications to which DPUs and SmartNICs are being applied. Additionally, we have separated out those works connecting these devices with HPC, as well as various forms of machine learning (ML) and artificial intelligence (AI) applications, as separate areas of interest.
@@ 6 Application Areas of DPUs
Since the line is often very blurred between infrastructure and more unconventional applications of SmartNICs, in this section we will also present those works which experiment with using SmartNICs for various other kinds of offloads and co-processing tasks, reaching far into the realm of applications as well. Although they are quickly becoming crucial in infrastructure acceleration, the flexibility of DPUs has led to their skyrocketing adoption in a surprising range of domains beyond that niche. From blockchain technologies to molecular dynamics simulations, they are being treated as co-processors, data management and analytics devices, AI-accelerators, and more. Table 6 gives a reasonable breakdown of the many overlapping topics, and their sources, within this category.
@@
Hardware Implementation, FPGAs [84, 70, 100, 83] (and Section 6.3)
Distributed Storage [84, 20, 17, 99]
@@
6.1 DPUs in Commercial and Infrastructure Applications
The commercial application of DPUs has become increasingly prevalent in modern computing environments. As they are primarily infrastructure devices, much of the current research around DPUs is in using them to improve the infrastructure itself of the data centers they belong to; their ability to accelerate network communication and offload tasks relating to packet processing, security, infrastructure processing, storage, virtualization, cloud computing, etc. can provide an important advantage to the companies that use them. Since these are the primary familiar and intended applications of SmartNICs, these areas of research tend to show many positive results, as described in the papers we address in this section.
[…]
A potential way to use SmartNICs to improve disk and network I/O operations is by rearchitecting the TCP stack itself, as done in [71] by Kim et al. They present ‘IO-TCP’, which improves content delivery by splitting its stack of operations between the host CPU and the SmartNIC. Since offloading the whole stack to the SmartNIC would leave data inaccessible to the CPU, and the SmartNIC is not fast enough to compete for some operations, they differentiate between “control plane” operations to be handled by the CPU, and “data plane” operations to be managed by the SmartNIC. This reduces cache pollution between operations with large amounts of disk I/O. Their work is tested on the BlueField-2 DPU, demonstrating approximately double the throughput in cases with high CPU congestion, even compared to systems using far more CPU cores.
@@ 6.1 DPUs in Commercial and Infrastructure Applications
Hardware and FPGA-Based; Network, Packets, & Transactions: Some works, such as that of Zhang et al. in [84], aim to improve the design of SmartNICs themselves. Their efforts demonstrate ‘SmartDS’, an FPGA prototype SmartNIC with the ability to split message headers and bodies between the host and SmartDS, offloading compression and improving throughput for distributed storage systems, with potential for accelerating and reducing the number of middle-tier servers required for cloud storage systems, etc. Their tests show up to 4.3x throughput and a latency reduction of 2.6x. Similarly, Li et al. develop ‘AlNiCo’ [70], an on-NIC network transaction scheduling system which aims to offload the scheduling of transaction requests and responses to specific host CPUs, in a manner which reduces both overhead and resource contention, and which adapts to feedback during dynamic workloads. Their solution uses an FPGA-based SmartNIC to program their own hardware acceleration, into which they feed feature vectors comprised of request information. They also use RDMA (Remote Direct Memory Access), and the development of their own “scheduling-enabled” RPC (Remote Procedure Call). They show improvements to both throughput and latency. Also, Yizhou et al. [83] fill a niche by presenting what they claim is the first SmartNIC that is both multi-tenant for virtualization needs, and programmable but hardware-based, which they call ‘SuperNIC’ (in a naming collision with NVIDIA’s SuperNIC products). They focused on optimizing execution of entire connected groups (flows, or directed acyclic graphs) of network tasks, and included various optimization techniques resulting in low overhead and high throughput compared to the baseline. Furthermore, the researchers proposed a fair sharing mechanism for the SuperNIC’s hardware resources. As for traffic analysis, Siracusano et al. [98] use a hardware-implemented binary neural network, described in section 6.3), to speed things up by two orders of magnitude. Finally, Brunella et al.’s ‘Hyperion’ [100], described later, is a more direct example of modifying DPU hardware, in pursuit of better distributed storage systems.
@@
Distributed File Systems: Others have also been able to achieve substantial speedups by focusing on distributed file systems (DFS). LineFS, a persistent memory (PM) distributed file system (DFS), was first presented by Jongyul Kim et al. in a paper describing its design, implementation, and evaluation [88]. LineFS resides within the SmartNIC itself to offload DFS operations and reduce CPU contention, while still taking advantage of new client-local persistent memory (PM) to create speedups by preferring client-local behavior. They test their work on a BlueField and compare it to Assise (a current PM DFS), showing improvement in DFS availability, latency of LevelDB operations by up to 80%, and throughput in Filebench by as much as 79%. However, [23] notes a limited range of situations in which LineFS improves performance. Wei et al. [81] perform case studies using both LineFS (an on-SmartNIC DFS) and DrTM-KV (an RDMA-based disaggregated KV-store), focusing on the management of such systems, and the idea of using multiple communication paths concurrently, achieving up to 30% and 25% improvements, respectively, using a BlueField-2 DPU.
Gootzen et al. [17] propose ‘DPFS’, which represents a decoupling of the host from its file system and using virtio-fs to virtualize it onto the DPU. The host’s computation is freed up significantly, with little latency cost for tenants. K. Zhong et al. likewise proposed ‘DPC’ [85], a high-performance, DPU-accelerated, distributed/standalone file system client, designed to offload computationally intensive tasks to DPUs. DPC optimizes file operations through ‘nvme-fs’, an enhanced NVMe protocol that enables low-latency and high-performance interactions. Additionally, they introduce ‘KVFS’, a key-value-based file system that replaces underutilized local disks with disaggregated storage. Experimental results show significant improvements, including over 80% CPU usage savings in high-concurrency scenarios, and a 90% CPU usage reduction in distributed file systems.
@@
The conceptual design itself of a DPU is altered by Brunella et al.’s ‘Hyperion’, which replaces host CPU-based devices as the primary manager of distributed NVMe storage [100], reimagining typical architectures to eliminate the need for many CPU server devices.
@@
Compression: As some DPUs contain compression ASICs, it is demonstrable that compression time and latency can be drastically reduced via their use. Li et al. [72] introduced a library called ‘PEDAL’ that optimizes data compression designs utilizing the DPU’s hardware capabilities. They also highlight the challenges of existing compression methods due to high computational demands, and present a co-design with the MPICH MPI library which achieves a speedup of up to 101x in compression time, also reducing the communication latency by up to 88x, compared to not using ASICs. Lastly, they focused on enhancing the data compression in communication-oriented HPC scenarios. Ding et al. [86, 87] also choose to offload compression to the DPU with ‘D2Comp’, in order to accelerate LSM-tree compaction; Ravi et al. [47] do so with ‘Runway’ for in-transit computation; and Zhang et al. [84] with SmartDS (all further detailed in other parts of this survey). Others, such as Jianshen [18], make use of the compression ASICs as well.
@@
KV-Stores, Databases, & Big Data Systems: SmartNICs are being used to offload various aspects of distributed data stores, with many claims that these offloads appear to be effective. In [91], Lin et al. focus their efforts on the shuffle processes (map-reduce, etc.) of data-intensive applications, and more generally on dynamically partially offloading operations requiring intermediate data exchange between compute nodes. They present ‘SmartShuffle’ for coordinating offload to maximize usage of both the CPU and SmartNIC simultaneously. They attach it to Spark and test it on a Broadcom Stingray DPU, achieving up to 40% faster benchmark performance than Spark. Somewhat related, Schuh et al. [37] introduce ‘Xenic’, a system that employs an asynchronous, aggregated execution paradigm with adaptable point-to-point communication between SmartNICs, improving both network and core efficiency in performing sharded data store transactions. Their experimental findings demonstrate that Xenic achieves notable improvements on three different benchmarks, at least doubling throughput and reducing latency. However, [23] questions Xenic’s use in off-path SmartNICs.
@@
As for more specific aspects of such systems, Ding et al. [86, 87] proposed another improvement, ‘DComp’ and then ‘D2Comp’, wherein they integrate DPU-offloaded LSM-tree compaction with RocksDB. This approach was evaluated using on a BlueField-2, accelerating compaction performance by up to 4x, write and read throughput somewhat, and reduce host CPU contention. Also, Shi et al. [73] proposed a set of coherent in-network, tripartite graph-based erasure coding (EC) primitives, named ‘INEC’ (building on their previous work named ‘TriEC’ [103]), to more easily allow offloading of EC to SmartNICs despite the variety of schemes EC follows. The researchers demonstrated that NIC implementations of INEC primitives, for various EC schemes, with a co-designed KV store, results in improved throughput and write performance.
@@
Lastly, when dealing with network data flows through SmartNICs, Liu et al. [18] recommend the use of Apache Arrow at the foundation, because its data format can save data transformation time. The authors share their experience in adjusting a partitioning algorithm for particle data to work with Apache Arrow.
Works belonging to other sections, such as that by Wei et al. [81], may also deal with KV-stores.
@@
5G Network Infrastructure: Yan et al. [60] introduced a SmartNIC solution that is P4-enabled and FPGA-based. This solution was specifically developed to cater to the networking requirements of web-scale cloud and the 5G / beyond-5G era. The authors demonstrated the application of this solution in a 5G environment, with a particular focus on network slicing from edge data center to core data center. Their experiments achieve a throughput of up to 84.8 Gbps, even using only a single CPU core, and up to 30% higher bandwidth. Later, Borromeo et al. [90] add a 5G DU Low-PHY layer to an FPGA SmartNIC using the OpenCL framework. They demonstrate that incorporating these 5G functionalities into a SmartNIC as an offload, as opposed to other current solutions for a 5G network, can result in notable reductions in both processing time and power consumption.